# MUSCLE: A Model Update Strategy for Compatible LLM Evolution

### v1

@@ A.1 Training Hyperparameters and Evaluation
We show an overview of the design choices of our LoRA adapter for task and compatibility training in Table 7. We learn on the entire context and answer tokens during fine-tuning. We find that successfully training the compatibility adapter generally requires larger LoRA rank. Task adapters can be trained with lower rank, but also experience performance boost with higher rank. We use the training splits of the respective datasets where we use 80% of the training set as training data, and 20% as validation data for selecting the best model based on validation loss. We evaluate for selection after each epoch. All compatibility adapter models are initialized with the previously trained task adapter models for version 2. KL divergence temperature is set to 2. We use deepspeed for optimization Rasley et al. (2020).
For benchmark evaluation, we use LM evaluation harness Gao et al. (2023), where most of the benchmarks (GSM8k, HellaSwag, PIQA) are already defined, and we use them as-is. We add evaluation for SAMsum summarization evaluating ROUGE-1 score with a no repeat N-gram size of 2, and generation stop words of “Dialogue:”, “Summary:” (which are the key words for the context and answer behavior), “</s>” and double new lines. All of our tasks are based on the English language. All Experiments were run on NVIDIA A100 and H100 GPUs with a compute budget of 720x8 GPUh.
A.2 Model Update Evaluation
In the evaluation of regression for different model updates (Fig. 4), we consider the model pairs shown in Table 8.
@@ A.2 Model Update Evaluation
Table 8: Model pairs used to analyse regression in model updates.
A.3 Downstream Tasks
@@
We describe the datasets used in our experiments. Each dataset poses unique challenges and targets different aspects of language understanding and reasoning.
@@
The Grade School Math 8K (GSM8K) dataset designed to evaluate mathematical reasoning capabilities. It consists of grade school-level math word problems. These problems require the application of mathematical concepts and the ability to reason about quantities and operations in a textual format. The dataset is structured to test the performance of models on arithmetic, algebra, geometry, and statistics problems, reflecting a wide range of mathematical knowledge in education. We use this dataset as a representative case to test problem solving cases. We evaluate regression in this task with exact match, which is a strict evaluations on a fraction of produced tokens, but does not account for the reasoning process.
@@
The SAMSum dataset consists of dialogue summaries designed to facilitate the evaluation of automatic conversational summarization models. It contains dialogue instances, paired with human-written summaries. These conversations mimic real-life scenarios to enable learning generate coherent and concise summaries. We use this dataset as a representative case to test language generation behavior. We evaluate regression in this task with ROUGE-1, which is a lenient evaluations on a fraction of produced tokens that are contained in the ground truth data.
@@
HellaSWAG is a dataset for assessing common sense reasoning and predictive text modeling. It builds on the SWAGZellers et al. (2018) dataset by providing more challenging distractors. HellaSWAG consists of multiple-choice scenarios where a model must predict the most plausible continuation among four options, focusing on everyday activities, scenarios, and interactions. We use this dataset as a representative case to test abilities that require not just linguistic understanding but also real-world knowledge and common sense reasoning. We evaluate regression in this task with log-likelihoods for the correct answer, as multiple choices are given.
@@
The Physical Interaction QA (PIQA) dataset tests the understanding of physical and causal interactions in the real world through textual descriptions. It contains scenarios that require reasoning about physical properties, actions, and outcomes. Each scenario is presented as a question with two possible solutions, where the model must choose the most physically plausible one. We use this dataset as a representative case for evaluating models on tasks that require an understanding of the physical world and its governing principles. We evaluate regression in this task with log-likelihoods for the correct answer, as different choices are given.

### v2

@@ A.1 Training Hyperparameters and Evaluation
We show an overview of the design choices of our LoRA adapter for task and compatibility training in Table 8. We learn about the entire context and answer tokens during fine-tuning. We find that successfully training the compatibility adapter generally requires a larger LoRA rank. Task adapters can be trained with lower rank, but also experience performance boost with higher rank. We use the training splits of the respective datasets where we use 80% of the training set as training data, and 20% as validation data for selecting the best model based on validation loss. We evaluate for selection after each epoch. All compatibility adapter models are initialized with the previously trained task adapter models for version 2 model. KL divergence temperature is set to 2. We use deepspeed for optimization Rasley et al. (2020).
@@
For benchmark evaluation, we use LM evaluation harness Gao et al. (2023), where most of the benchmarks (GSM8k, HellaSwag, PIQA) are already defined, and we use them as-is. We add an evaluation for SAMsum summarization evaluating ROUGE-1 score with a no-repeat N-gram size of 2, and generation stop words of “Dialogue:”, “Summary:” (which are the keywords for the context and answer behavior), “</s>” and double new lines. All of our tasks are based on the English language.
A.2 Training Cost
All Experiments were run on NVIDIA A100 and H100 GPUs with an overall compute budget of 720x8 GPUh. Assuming hourly rates of 2.5\$, this would amount to 14,400\$. Compared to regular parameter-efficient fine-tuning, our compatibility method requires both model versions in GPU memory during training, hence fewer data batches can be processed per time step. However, as the compatibility adapters are initialized with the task adapters, we can view it as a continued training. There are no differences for inference time and costs compared to regular parameter-efficiently fine-tuned LoRa adapters.
A.3 Model Update Evaluation
In the evaluation of model update regression for different model updates (Fig. 4), we consider the model pairs shown in Table 9.
@@ A.3 Model Update Evaluation
Table 9: Model pairs used to analyse model update regression.
A.4 Auxiliary Cross-Entropy Loss
For model updates with large performance gaps between $\mathcal{M}_{v1}$ and $\mathcal{M}_{v2}$, we find that an auxiliary cross-entropy loss enhances the stability of training. When the performance of $\mathcal{M}_{v1}$ is greatly inferior to $\mathcal{M}_{v2}$, we assume that the inferior $\mathcal{M}_{v1}$ introduces errors that steer the model away too much from the ground truth. We can account for this by adding the cross-entropy loss to align with the ground truth. Using Eq. 1, we add a $\mathcal{L}_{CE}$ for training (scaled with hyperparameter $\lambda$) .
$\mathcal{L}_{CE}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{K}y_{i,k}\log(\sigma(z_{\mathcal{M}^{C}_{v2},i,k}))$ (4)
$\mathcal{L}=\lambda\mathcal{L}_{comp}^{m}+(1-\lambda)\mathcal{L}_{CE}$ (5)
We find that likelihood-based masking (Eq. 2) also requires auxiliary cross-entropy loss for stability. Even though one model version might provide a higher likelihood than the other, this likelihood does not necessarily mean a suitable probability distribution over the vocabulary, hence adding alignment to the ground truth helps to provide better results.
A.5 Downstream Tasks
@@
We describe the datasets used in our experiments. Each dataset poses unique challenges and targets different aspects of language understanding and reasoning. Dataset sizes are shown in Table 10.
@@
The Grade School Math 8K (GSM8K) dataset is designed to evaluate mathematical reasoning capabilities. It consists of grade school-level math word problems. These problems require the application of mathematical concepts and the ability to reason about quantities and operations in a textual format. The dataset is structured to test the performance of models on arithmetic, algebra, geometry, and statistics problems, reflecting a wide range of mathematical knowledge in education. We use this dataset as a representative case to test problem-solving cases. We evaluate model update regression in this task with exact match, which is a strict evaluation on a fraction of produced tokens, but does not account for the reasoning process.
@@
The SAMSum dataset consists of dialogue summaries designed to facilitate the evaluation of automatic conversational summarization models. It contains dialogue instances, paired with human-written summaries. These conversations mimic real-life scenarios to enable learning to generate coherent and concise summaries. We use this dataset as a representative case to test language generation behavior. We evaluate model update regression in this task with ROUGE-1.
@@
HellaSWAG is a dataset for assessing common sense reasoning and predictive text modeling. It builds on the SWAG Zellers et al. (2018) dataset by providing more challenging distractors. HellaSWAG consists of multiple-choice scenarios where a model must predict the most plausible continuation among four options, focusing on everyday activities, scenarios, and interactions. We use this dataset as a representative case to test abilities that require not just linguistic understanding but also real-world knowledge and common sense reasoning. We evaluate model update regression in this task with log-likelihoods for the correct answer, as multiple choices are given.
@@
The Physical Interaction Question Answering (PIQA) dataset tests the understanding of physical and causal interactions in the real world through textual descriptions. It contains scenarios that require reasoning about physical properties, actions, and outcomes. Each scenario is presented as a question with two possible solutions, where the model must choose the most physically plausible one. We use this dataset as a representative case for evaluating models on tasks that require an understanding of the physical world and its governing principles. We evaluate model update regression in this task with log-likelihoods for the correct answer, as different choices are given.